OpenStudy (anonymous):

Solve the following system. x − y = 10 x + y = 2 (6, −4) (6, 4) (−4, 6) (−6, 4) please this is kinda of confusing me here i need help

OpenStudy (anonymous):

Argh this is confusing me too, give me a second.

OpenStudy (anonymous):

Subtract each other x-y=10 x+y=2 -2y=8 y=-4 Then put y back into one of the equations to solve for X x-(-4)=10 x=6 or.. x+(-4)=2 x=6 So the answer is (6,-4)